OpenStudy (anonymous):

John has a jar filled with 150 coins made up of quarters and nickels. The total value of the coins is $26.70. How many quarters and how many nickels are in the jar? A. 75 quarters; 75 nickels B. 96 quarters; 54 nickels C. 104 quarters; 46 nickels D. 54 quarters; 96 nickels

OpenStudy (campbell_st):

well its a simultaneous equations question q + n = 150 0.25q + 0.05n = 26.70 now solve them simultaneously

OpenStudy (anonymous):

that looks like an explanation to me.. ^ haha people do teach on this website nice to know

OpenStudy (campbell_st):

thanks

OpenStudy (anonymous):

lol yw I was just proving my point to an ignorant person

OpenStudy (anonymous):

Can you tell me the answer please? I'm just trying to finish the last questions of my math homework so I can get some sleep.

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

that doesn't make your point right dummy and I explain

OpenStudy (anonymous):

It's finals week and I just want some sleep. Thank you for giving me a quick answer Shay, and thank you for taking your time to explain @campbell_st

OpenStudy (anonymous):

@campbell_st explained it why repeat what he said all you're doing is substituting in the amount of nickles and quarters

OpenStudy (anonymous):

It's pretty obvious looking at the formula

OpenStudy (anonymous):

@campbell_st this is the ignorance I was referring to :/

OpenStudy (anonymous):

B is the answer, just plug in

OpenStudy (anonymous):

Shay was correct. You other guys take things way too seriously.

OpenStudy (campbell_st):

I can tell you how to solve the equation.... multiply the 2nd equation by -4 and then add the 2 equations... this will eliminate the quarters (q) and allow you to find how many nickels you have so you get q + n = 150.00 -q - 0.2n = -106.80 ---------------------- so add the 2 equations and what do you get...?

OpenStudy (anonymous):

Thank you @kk54 we don't mind the haters

OpenStudy (anonymous):

I never said I was mad and your just annoying af like a d.amn fly that just won't go away

OpenStudy (anonymous):

You really have no life .. You don't see how childish you are you need to grow up seriously

OpenStudy (anonymous):

Anyways @kk54 anymore questions feel free to let me know

OpenStudy (anonymous):

Haha why you delete your comments ignorant person ? You had some time to reflect on your bad behavior ?